# New entropy, thermodynamics of apparent horizon and cosmology

### v3

@@ 4.1 Inflation
Inflation corresponds to the accelerated expansion of the universe and gravity possesses a repulsive force, At the early stage of the universe, i.e. inflation, the matter is absent, $\rho=p=0$. The entropic energy (dark energy) density triggers the universe to accelerate. Then we obtain from Eq. (16) that $\dot{H}=0$, i.e. during inflation the Hubble parameter is a constant. As a result, the scale factor is given by $a(t)=a_{0}\exp(Ht)$ and corresponds to de Sitter spacetime. It should be noted that de Sitter stage corresponds to the eternal inflation. The Hubble parameter at the early universe (inflation) is $H\approx 10^{-3}M_{Pl}$ ($M_{Pl}=1/\sqrt{16\pi G}$). Then $GH^{2}\approx 2\cdot 10^{-8}$. From Eq. (17), at $\rho=0$ we obtain the equation as follows:
$H^{2}-\frac{b^{2}}{H^{2}+b}-2b\ln\left(\frac{H^{2}+b}{b}\right)=0.$ (32)
Introducing new variable $x=H^{2}/b$, Eq. (32) becomes
$x-\frac{1}{x+1}-2\ln(x+1)=0.$ (33)
The non-trivial solution to Eq. (33) is given by $x=3.04002$ and we have the relation
@@
Knowing the Hubble parameter during universe inflation, $H\approx 10^{-3}M_{Pl}$, we can fix parameter $b$ from Eq. (34) and the entropy parameter $\gamma=bG/\pi$. Thus, we obtain $\gamma\approx 2\times 10^{-9}$. It is worth noting that to have a viable model of inflation one needs a smooth graceful exit into the deceleration stage [1].
@@ 5 F(T)-gravity from generalized entropy
In teleparallel gravity, a scalar torsion $T$ plays a role of a fundamental field similar to the curvature $R$ in Einstein’s general relativity theory [42, 43]. To describe the inflationary era and the current universe accelerating expansion, one introduces a Lagrangian in the form of $F(T)$ in analogy with $F(R)$-gravity. The torsion scalar field $T$ is given by [44, 45]
@@
$\left[F(T)-2TF^{\prime}(T)\right]=-T-\frac{36b^{2}}{6b-T}-12b\ln\left(1-\frac{T^{2}}{6b}\right).$ (38)
@@
$F(T)=T-3\sqrt{6bT}~{}\mbox{tanh}^{-1}\sqrt{\frac{T}{6b}}$
$-12b\ln\left(1-\frac{T}{6b}\right)+C\sqrt{T}-6b,$ (39)
where $C$ is the integration constant. Thus, the teleparallel gravity with the function (39) corresponds to entropic cosmology with entropy (15) under consideration.
@@ 6 Conclusion
Considering a novel entropy $S_{K}=S_{BH}/(1+\gamma S_{BH})$ we have a property similar to the Bekenstein–Hawking entropy $S_{BH}$. By using our entropy and by employing the holographic principle, we have obtained a dark energy model. The $S_{K}$ vanishes when the apparent horizon radius $R_{h}$ is zero and it increases monotonically when the apparent horizon radius $R_{h}$ increases. The barotropic perfect fluid and spatial flat FLRW universe were studied. We have obtained, from first law of apparent horizon thermodynamics, the modified Friedmann’s equations. The second Friedmann’s equation possesses an addition term corresponding to the density of dark energy which can be treated as a dynamical cosmological constant. Implying that there is no interaction between various components of cosmos, and the dark energy density and pressure $\rho_{D}$ and $p_{D}$ obeys ordinary conservation law, we computed $\rho_{D}$ and $p_{D}$. The EoS parameter for dark energy $w_{D}=p_{D}/\rho_{D}$ was calculated and it was shown that $\lim_{H\rightarrow\infty}w_{D}=-1$. Thus, at the small apparent horizon radius $R_{h}$ the de Sitter stage is realized which describes the inflation. In the model under consideration the universe can have two phases, acceleration and deceleration, that is due to holographic dark energy. It should be noted that Barrow and Tsallis entropies also lead to cosmology which ia due to Einstein’s equations with the dynamical cosmological constant [40]. We computed the deceleration parameter that shows the possibility, for some model parameters, to describe the acceleration at the current era. The transition redshifts (when $q=0$) was calculated (given in Table I) for some parameters. We show (see Table II) that at $b\approx 0.1$ and $w=-2/3$ we have the deceleration parameter $q_{0}\approx-0.535$ which was observed in universe at the current era [39]. Thus, our approach, based on new entropy and leading to modified Friedmann equations, can describe universe inflation and late time universe acceleration. It has been proven that entropic cosmology with our entropy proposed is equivalent to cosmology based on the teleparallel gravity with the function $F(T)$.

### v4

@@ 4.1 Inflation
Inflation corresponds to the accelerated expansion of the universe and gravity possesses a repulsive force. At the early stage of the universe, i.e. inflation, the matter is absent, $\rho=p=0$. The entropic energy (dark energy) density triggers the universe to accelerate. Then we obtain from Eq. (16) that $\dot{H}=0$, i.e. during inflation the Hubble parameter is a constant. As a result, the scale factor is given by $a(t)=a_{0}\exp(Ht)$ and corresponds to the de Sitter spacetime. It should be noted that de Sitter stage leads to the eternal inflation. The Hubble parameter at the early universe (inflation) is $H\approx 10^{-3}M_{Pl}$ ($M_{Pl}=1/\sqrt{8\pi G}$). Then $GH^{2}\approx 4\cdot 10^{-8}$. From Eq. (17), at $\rho=0$ and using the variable $x=H^{2}/b$ we obtain the equation as follows:
$x-\frac{1}{x+1}-2\ln(x+1)=0,$ (40)
which is identical with the Eq. (29) for the phantom divide. The non-trivial solution to Eq. (40) is given by $x=3.04002$ and we have the relation
@@
Knowing the Hubble parameter during universe inflation, $H\approx 10^{-3}M_{Pl}$, we can fix parameter $b$ from Eq. (41) and the entropy parameter $\gamma=bG/\pi$. Thus, we obtain $\gamma\approx 4.2\times 10^{-9}$. It is worth noting that to have a viable model of inflation one needs a smooth graceful exit into the deceleration phase [1].
@@ 5 F(T)-gravity from generalized entropy
In teleparallel gravity, a scalar torsion $T$ plays a role of a fundamental field similar to the curvature $R$ in Einstein’s general relativity theory [43, 44]. To describe the inflationary era and the current universe accelerating expansion, one introduces a Lagrangian in the form of $F(T)$ in analogy with $F(R)$-gravity. The teleparallel theory of gravity based on teleparallel geometry uses the Weitzenböck connection (not the Levi-Civita connection). In the $F(T)$ theory the field equations are the second order which is an advantage compared with $F(R)$ theory having the fourth order equations. The spacetime does not have the curvature and possesses the torsion. The torsion scalar field $T$ is given by [45, 46]
@@
$F(T)-2TF^{\prime}(T)=-T-\frac{36b^{2}}{6b-T}-12b\ln\left(1-\frac{T^{2}}{6b}\right).$ (45)
@@
$F(T)=T-3\sqrt{6bT}~\mbox{tanh}^{-1}\sqrt{\frac{T}{6b}}-12b\ln\left(1-\frac{T}{6b}\right)+C\sqrt{T}-6b,$ (46)
where $C$ is the integration constant. It is worth noting that the term $C\sqrt{T}$ in Eq. (45) is the solution of the homogeneous equation $F(T)-2TF^{\prime}(T)=0$ and should be eliminated, $C=0$. As $T=-6H^{2}<0$, we can use the formula $i~\mbox{tanh}^{-1}(ix)=-\arctan(x)$. Then Eq. (46) is converted into the real function (at $C=0$) as follows:
$F(T)=T+3\sqrt{-6bT}~\arctan\sqrt{-\frac{T}{6b}}-12b\ln\left(1-\frac{T}{6b}\right)-6b.$ (47)
Some models of teleparallel gravity with different functions $F(T)$ were studied in [48, 49] Thus, the teleparallel gravity with the function (47) corresponds to entropic cosmology with entropy (15) under consideration.
@@ 6 Conclusion
Considering a novel entropy $S_{K}=S_{BH}/(1+\gamma S_{BH})$ we have a property similar to the Bekenstein–Hawking entropy $S_{BH}$. By using our entropy and by employing the holographic principle, we have obtained a dark energy model. The $S_{K}$ vanishes when the apparent horizon radius $R_{h}$ is zero and it increases monotonically when the apparent horizon radius $R_{h}$ increases. The barotropic perfect fluid and spatial flat FLRW universe were studied. We have obtained, from first law of apparent horizon thermodynamics, the modified Friedmann equations. The second Friedmann equation possesses an addition term corresponding to the density of dark energy which can be treated as a dynamical cosmological constant. Implying that there is no interaction between various components of cosmos, and the dark energy density and pressure $\rho_{D}$ and $p_{D}$ obeys ordinary conservation law, we computed $\rho_{D}$ and $p_{D}$. The EoS parameter for dark energy $w_{D}=p_{D}/\rho_{D}$ was calculated and it was shown that $\lim_{H\rightarrow\infty}w_{D}=-1$. Thus, at the small apparent horizon radius $R_{h}$ the de Sitter stage is realized which describes the inflation. In the model under consideration the universe can have two phases, acceleration and deceleration, that is due to holographic dark energy. It should be noted that Barrow and Tsallis entropies also lead to cosmology which is due to Einstein’s equations with the dynamical cosmological constant [41]. We computed the deceleration parameter that shows the possibility, for some model parameters, to describe the acceleration at the current era. We show that at the entropy parameter $\gamma\approx 0.055GH_{0}^{2}$ and w=0.283w=-0.283 we have the deceleration parameter $q_{0}\approx-0.535$ and the normalized density parameter $\Omega_{m0}\approx 0.315$ which were observed at the current era [40]. We showed that our model possesses the phantom divide for the EoS of dark energy. Thus, our approach, based on new entropy and leading to modified Friedmann equations, can describe the universe inflation and the late time of universe acceleration. It has been proven that entropic cosmology with our entropy proposed is equivalent to cosmology based on the teleparallel gravity with the function $F(T)$ (Eq. (47)).